# Game-Time: Evaluating Temporal Dynamics in Spoken Language Models

### v1

@@ 1 Introduction
In the pursuit of human-like conversation with machines, the research frontier is moving beyond text-based Large Language Models (LLMs). The next challenge lies in mastering conversational dynamics in real-time speech, which has given rise to the field of conversational Spoken Language Models (SLMs) [1, 2, 3, 4, 5, 6, 7]. This marks a critical shift from rigid turn-by-turn dialogues to fluid spoken interactions. Achieving such dynamics requires SLMs to operate in a real-time full-duplex manner [8, 9], where models must listen and speak simultaneously while producing seamless responses. This is inherently difficult, demanding synchronous speech generation, continual intent recognition, and precise control over both what to respond and when to respond. A further challenge lies in modeling the temporal dynamics of spoken interaction, where contemporary systems often fail to capture the fine-grained timing that is essential for advanced conversational fluency. For instance, they struggle to process user speech while planning a coherent reply that aligns with user-specified timing and tempo. This limitation underscores a fundamental deficiency: the lack of time-awareness.
@@ 2.1 Full-duplex Spoken Language Models
(1) Dual-channel SLMs [22, 17, 27, 28] use two input channels: a listening channel for user speech and a speaking channel for the model’s own output. At each time step, the model simultaneously processes both inputs and generates output speech, which may be either voiced or explicitly silent. Although this architecture significantly increases modeling complexity, it naturally supports real-time listening and speaking concurrently.
(2) Time-multiplexing SLMs [29, 19, 18] include a state prediction mechanism [30] that decides whether to speak or remain silent. During user speech, the model withholds output and monitors for an appropriate moment to take the turn and respond. During its own speech, the model generates in an autoregressive manner, and this generation is often interrupted by incoming user speech.
@@ 2.2 Benchmarks for Conversational Spoken Language Models
A number of benchmarks have been proposed to evaluate SLMs. Some focus on the spoken language understanding and paralinguistics generation [10, 31, 32, 33, 34, 35]. Others, including Full-Duplex-Bench [12], Talking Turns [13], and FD-Bench [36] primarily assess dialogue behaviors including latency, back-channeling, and turn-taking, with an emphasis on the conversational naturalness. However, these benchmarks do not directly test a model’s ability to comprehend and act upon explicit temporal requirements. Game-Time complements these existing works by shifting the focus from what to say to when to say, introducing a framework for evaluating timing, tempo, and simultaneous speaking within a conversation.

### v2

@@ 1 Introduction
In the pursuit of human-like conversation with machines, the research frontier is moving beyond text-based Large Language Models (LLMs). The next challenge lies in mastering conversational dynamics in real-time speech, which has given rise to the field of conversational Spoken Language Models (SLMs) [17, 10, 2, 33, 18, 16, 13]. This marks a critical shift from rigid turn-by-turn dialogues to fluid spoken interactions. Achieving such dynamics requires SLMs to operate in a real-time full-duplex manner [22, 6], where models must listen and speak simultaneously while producing seamless responses. This is inherently difficult, demanding synchronous speech generation, continual intent recognition, and precise control over both what to respond and when to respond. A further challenge lies in modeling the temporal dynamics of spoken interaction, where contemporary systems often fail to capture the fine-grained timing that is essential for advanced conversational fluency. For instance, they struggle to process user speech while planning a coherent reply that aligns with user-specified timing and tempo. This limitation underscores a fundamental deficiency: the lack of time-awareness, which is essential for real-world deployment, enabling SLMs to deliver time-critical instructions (e.g., emergency guidance) and to synchronize speech with user activities (e.g., healthcare voice agents).
@@ 2.1 Full-duplex Spoken Language Models
(1) Dual-channel SLMs [23, 11, 29, 32] directly process two channels: a listening channel for user speech and a speaking channel for the model’s own output. Although this architecture significantly increases modeling complexity, it naturally supports real-time listening and speaking concurrently.
(2) Time-multiplexing SLMs [41, 30, 38] include a state prediction mechanism [28] that decides whether to speak or remain silent. The model monitors for turn-taking cues during user speech and switches to speech generation until external interruption occurs.
@@ 2.2 Benchmarks for Conversational Spoken Language Models
A number of benchmarks have been proposed to evaluate SLMs. Some focus on the spoken language understanding and paralinguistics generation [35, 1, 7, 14, 21, 39]. Others target interaction naturalness, including Full-Duplex-Bench [20], which evaluates SLM’s conversational abilities such as interruption handling and overlap management; Talking Turns [3], which focuses on turn-taking dynamics and conversational flow in dialogue; and FD-Bench [25], which assesses the SLM’s responsiveness and robustness under noisy scenarios. More recently, Beyond Words [19] proposes a multimodal framework to determine when a model should speak by modeling turn-taking and backchanneling in audiovisual conversations. While these benchmarks characterize important conversational behaviors, they do not directly evaluate whether models can comprehend and satisfy explicit temporal constraints.